OpenStudy (purplemexican):

Which equation is y = 3(x – 2)^2 – (x – 5)^2 rewritten in vertex form?

OpenStudy (anonymous):

do you know the general vertex form?

OpenStudy (purplemexican):

kinda

OpenStudy (anonymous):

lol okay well here

OpenStudy (anonymous):

\[y=3\left( x^2-4x+4 \right)-\left( x^2-10x+25 \right)\] \[=2x^2-2x-13=2\left( x^2-x+\frac{ 1 }{ 4 }-\frac{ 1 }{ 4 } \right)-13\] \[=2\left( x-\frac{ 1 }{ 2 } \right)^2-\frac{ 1 }{ 2 }-13=?\]

OpenStudy (purplemexican):